# Quantum-Corrected Thermodynamics of Conformal Weyl Gravity Black Holes: GUP Effects and Phase Transitions

### v1

@@ III Hawking Radiation of CWGBHs via Hamilton-Jacobi Formalism of Tunneling Particles
This expression represents a significant generalization of the standard Schwarzschild result and demonstrates the rich thermal structure emerging from CWG. The temperature formula is consistent with results obtained through alternative methods such as the null geodesic approach, providing an important cross-validation of the thermal spectrum derivation.
$r_{h}=1$ $r_{h}=5$ $r_{h}=10$
$k$ $\gamma$ $\beta$ $T_{H}$ $\Delta T_{H}$ $k$ $\gamma$ $\beta$ $T_{H}$ $\Delta T_{H}$ $k$ $\gamma$ $\beta$ $T_{H}$ $\Delta T_{H}$
0 0 0.1 0.0159 0 0 0.1 0.00127 0 0 0.1 0.000318
0 0 0.5 0.0796 +400.6% 0 0 0.5 0.00637 +401.6% 0 0 0.5 0.00159 +400.0%
0 0 1.0 0.1592 +901.3% 0 0 1.0 0.0127 +900.0% 0 0 1.0 0.00318 +900.0%
0 1 0.1 0.0963 +505.7% 0 1 0.1 0.0817 +6334% 0 1 0.1 0.0806 +25257%
0 1 0.5 0.1599 +905.7% 0 1 0.5 0.0868 +6733% 0 1 0.5 0.0819 +25677%
0 1 1.0 0.2396 +1406% 0 1 1.0 0.0932 +7240% 0 1 1.0 0.0836 +26195%
1 0 0.1 0.1759 +1006% 1 0 0.1 0.08127 +6298% 1 0 0.1 0.1593 +49906%
1 0 0.5 0.2396 +1406% 1 0 0.5 0.08637 +6699% 1 0 0.5 0.1606 +50315%
1 0 1.0 0.3192 +1907% 1 0 1.0 0.0927 +7200% 1 0 1.0 0.1623 +50849%
1 1 0.1 0.2563 +1512% 1 1 0.1 0.1627 +12717% 1 1 0.1 0.2399 +75377%
1 1 0.5 0.3199 +1912% 1 1 0.5 0.1678 +13118% 1 1 0.5 0.2412 +75786%
1 1 1.0 0.3996 +2412% 1 1 1.0 0.1742 +13622% 1 1 1.0 0.2429 +76321%
Table 1: Hawking temperature $T_{H}$ for CWGBHs across various parameter combinations of horizon radius $r_{h}$, cosmological parameter $k$, scale-dependent correction $\gamma$, and mass-like parameter $\beta$. The $\Delta T_{H}$ column shows percentage increases relative to the baseline case ($k=0$, $\gamma=0$, $\beta=0.1$) for each horizon radius, demonstrating the significant thermal enhancement effects introduced by CWG parameters.
Table 1 presents a comprehensive analysis of the Hawking temperature $T_{H}$ across different parameter regimes, revealing the complex interplay between the horizon radius $r_{h}$, the effective mass-like parameter $\beta$, the scale-dependent correction $\gamma$, and the cosmological analog $k$. The data demonstrates several key physical insights: For compact BHs with small horizon radii ($r_{h}=1$), the inverse square terms in Eq.(26) dominate the thermal behavior, making $T_{H}$ extremely sensitive to variations in both $\beta$ and $\gamma$ parameters. The percentage increases shown in the $\Delta T_{H}$ column illustrate that even modest changes in these parameters can lead to dramatic thermal enhancements, with temperature increases exceeding 2000% in some parameter combinations.
As the horizon radius increases to intermediate values ($r_{h}=5$) and large scales ($r_{h}=10$), the relative importance of the linear $kr_{h}$ term becomes increasingly pronounced, particularly when $k=1$. This transition reflects the changing balance between local gravitational effects (governed by $\beta$ and $\gamma$) and global spacetime geometry (controlled by $k$). The parameter $\beta$ consistently contributes to temperature enhancement across all scales, confirming its interpretation as a gravitational mass analog, while $\gamma$ introduces both direct constant contributions and nonlinear coupling effects with $\beta$, emphasizing its role in modifying the characteristic length scales of the theory.
IV GUP Corrected Hawking Radiation of CWGBHs
@@
To systematically account for the profound influence of quantum gravitational effects on BH thermodynamics, one of the most promising and well-developed theoretical approaches involves the implementation of the GUP. This principle introduces a fundamental minimal length scale that emerges naturally from several candidate theories of quantum gravity, including string theory, loop quantum gravity, and non-commutative geometry [76, 77, 78]. The GUP framework provides a phenomenological window into the quantum gravitational regime by modifying the standard Heisenberg uncertainty relations through the incorporation of Planck-scale corrections.
@@
$\Gamma_{\text{GUP}}\sim\exp\left(-\frac{2\pi E_{\text{GUP}}}{\kappa}\right)=\exp\left(-\frac{4\pi E}{B^{\prime}(r_{h})}\left(1+\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right)\right).$ (31)
@@
$T_{\text{GUP}}=\frac{B^{\prime}(r_{h})}{4\pi}\left(1+\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right)^{-1}=T_{H}\left(1+\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right)^{-1},$ (32)
where $T_{H}$ represents the classical Hawking temperature derived in the previous section. This result demonstrates that the emission spectrum undergoes a systematic redshift due to quantum gravitational corrections, with the magnitude of the effect being inversely related to the square of the horizon radius.
@@
V QC Thermodynamics of CWGBHs
@@
To provide a more complete theoretical framework, quantum statistical mechanics offers a sophisticated approach through systematic counting of horizon microstates [44, 45]. In canonical ensembles with fixed energy and particle number, quantum fluctuations give rise to subleading modifications to the classical entropy expressions. These corrections effectively capture the statistical uncertainties arising from quantum degrees of freedom that are entirely absent in semiclassical treatments, providing crucial insights into the quantum nature of gravitational thermodynamics [85, 86].
@@
[…]
where $T_{H}$ denotes the Hawking temperature derived in Eq.(26). Substituting the QC entropy expression and performing the integration yields
$E\approx\frac{1}{5}\pi k\,r_{h}^{5}+\frac{1}{8}\pi\gamma r_{h}^{4}-\frac{3}{4}\pi\,\beta^{2}\gamma r_{h}^{2}+\frac{1}{2}\pi\beta r_{h}^{2}.$ (35)
@@
Figure 1 presents a comprehensive density plot illustrating the behavior of the internal energy $E$ as a function of both the event horizon radius $r_{h}$ and the scale-dependent parameter $\gamma$, with fixed parameter values $\beta=2$ and $k=1$. The visualization clearly demonstrates that the energy increases monotonically with $r_{h}$, consistent with the thermodynamic interpretation of BH mass. The contour structure reveals that $\gamma$ exerts a negligible influence on $E$ at larger horizon scales, but becomes increasingly important in the small $r_{h}$ regime. This behavior indicates that scale-dependent geometric corrections subtly modify the BH energy structure near the Planckian regime, suggesting that quantum corrections encoded in $\gamma$ become critically relevant in the small-scale limit and may significantly influence evaporation dynamics or remnant formation scenarios.
Figure 1: Density plot of the inertial energy $E$ as a function of the event horizon radius $r_{h}$ and the scale-dependent parameter $\gamma$, with fixed values $\beta=2$ and $k=1$. The energy increases rapidly with $r_{h}$, while the influence of $\gamma$ is only significant in the small $r_{h}$ regime, highlighting the minimal but non-negligible role of scale-dependent corrections at short distances.
@@
The Helmholtz free energy, which provides crucial information about the system’s equilibrium properties, is defined through the standard thermodynamic relation
[…]
yielding upon substitution and integration
$F\approx-\frac{\pi k\,r^{5}}{20}-\frac{3\pi\,\beta^{2}\gamma r^{2}}{8}+\frac{\pi\beta r^{2}}{4}-\frac{kr}{2\pi}+\frac{3\beta^{2}\gamma}{4\pi r^{2}}-\frac{\beta}{2\pi r^{2}}.$ (37)
@@
Figure 2 displays the Helmholtz free energy behavior, revealing the complex interplay between geometric parameters and thermodynamic stability. The QC pressure, fundamental for understanding the mechanical properties of the BH system, follows from the thermodynamic relation [86]
@@
$P\approx\frac{r^{2}\pi k}{16\pi}+\frac{3\pi\,\beta^{2}\gamma}{16\pi r}-\frac{\pi\beta}{8\pi r}+\frac{k}{8\pi r^{2}\pi}+\frac{3\beta^{2}\gamma}{8\pi r^{5}\pi}-\frac{\beta}{4\pi r^{5}\pi}.$ (39)
Figure 2: Helmholtz free energy $F$ displaying the QC thermodynamic landscape for CWGBHs, showing the complex parameter dependence that governs equilibrium stability conditions.
Figure 3: QC pressure $P$ as a function of horizon parameters, illustrating the mechanical response properties of CWGBHs under varying geometric conditions.
@@
resulting in the detailed expression
$H\approx\frac{17\pi k\,r^{5}}{60}-\frac{\pi\,\beta^{2}\gamma r^{2}}{2}+\frac{\pi\gamma r^{4}}{8}+\frac{\pi\beta r^{2}}{3}\\ +\frac{kr}{6\pi}+\frac{\beta^{2}\gamma}{2\pi r^{2}}-\frac{\beta}{3\pi r^{2}}.$ (41)
Figure 4: Enthalpy $H$ distribution demonstrating the energy content of CWGBHs under constant pressure conditions, revealing the quantum-enhanced thermodynamic structure.
@@
$G\approx\frac{\pi k\,r^{5}}{30}-\frac{\pi\,\beta^{2}\gamma r^{2}}{8}+\frac{\pi\beta r^{2}}{12}-\frac{kr}{3\pi}+\frac{5\beta^{2}\gamma}{4\pi r^{2}}-\frac{5\beta}{6\pi r^{2}}.$ (42)
Figure 5: Gibbs free energy $G$ characterizing the global thermodynamic stability landscape and phase transition structure of QC CWGBHs.
@@
leading to the final result
$C\approx\frac{\pi^{2}\left(2k\,r^{3}-3\beta^{2}\gamma+\gamma r^{2}+2\beta\right)r^{4}}{k\,r^{3}+3\beta^{2}\gamma-2\beta}.$ (44)

### v2

@@ 3 Hawking Radiation of CWGBHs via Hamilton-Jacobi Formalism of Tunneling Particles
This expression represents a significant generalization of the standard Schwarzschild result and demonstrates the rich thermal structure emerging from CWG. The temperature formula is consistent with results obtained through alternative methods such as the null geodesic approach, providing an important cross-validation of the thermal spectrum derivation. It is important to note that for physically consistent BH solutions, the horizon radius $r_{h}$ must be determined self-consistently from the horizon condition $B(r_{h})=0$, which couples $r_{h}$ to the MK parameters $(\beta,\gamma,k)$. For a comprehensive analysis of the horizon structure in the MK metric, see Ref. [146].
$\beta=0.5$ $\beta=1.0$ $\beta=2.0$
$k$ $\gamma$ $r_{h}$ $T_{H}$ $\Delta T_{H}$ $k$ $\gamma$ $r_{h}$ $T_{H}$ $\Delta T_{H}$ $k$ $\gamma$ $r_{h}$ $T_{H}$ $\Delta T_{H}$
0 0 1.000 0.0796 0 0 2.000 0.0398 0 0 4.000 0.0199
0 0.05 0.988 0.0825 +3.6% 0 0.05 1.952 0.0426 +7.1% 0 0.05 3.817 0.0226 +13.4%
0 0.10 0.976 0.0852 +7.1% 0 0.10 1.908 0.0451 +13.4% 0 0.10 3.657 0.0246 +23.8%
0.05 0 0.956 0.0946 +18.9% 0.05 0 1.738 0.0665 +67.2% 0.05 0 2.847 0.0619 +211.3%
0.10 0 0.922 0.1083 +36.1% 0.10 0 1.595 0.0880 +121.1% 0.10 0 2.478 0.0913 +358.8%
0.05 0.05 0.946 0.0970 +21.9% 0.05 0.05 1.710 0.0679 +70.7% 0.05 0.05 2.777 0.0612 +207.4%
0.10 0.10 0.905 0.1123 +41.1% 0.10 0.10 1.551 0.0889 +123.3% 0.10 0.10 2.354 0.0856 +330.4%
Table 1: Hawking temperature $T_{H}$ for CWGBHs organized by mass parameter $\beta$. The horizon radii $r_{h}$ are self-consistently determined from $B(r_{h})=0$, ensuring physically realizable BH configurations. The baseline case for each $\beta$ corresponds to the Schwarzschild limit ($\gamma=0$, $k=0$), where $r_{h}=2\beta$. The percentage deviation $\Delta T_{H}$ quantifies the thermal enhancement relative to Schwarzschild at the same $\beta$.
[…]
4 GUP Corrected Hawking Radiation of CWGBHs
Before proceeding with the GUP analysis, we address a subtle but important theoretical point concerning the compatibility of the Generalized Uncertainty Principle with Conformal Weyl Gravity. The CWG action is constructed from the Weyl tensor squared and possesses local conformal symmetry, which forbids dimensionful coupling constants such as Newton’s constant $G$ in the fundamental Lagrangian. Since the GUP introduces a minimum measurable length scale $\Delta x_{\min}\sim\sqrt{\lambda}\,\ell_{P}$, where $\ell_{P}=\sqrt{\hbar G/c^{3}}$ is the Planck length, one might question whether these frameworks are mutually consistent.
The resolution lies in the well-established mechanism of spontaneous symmetry breaking. As demonstrated by Mannheim and Kazanas [85, 86], when conformally coupled scalar fields acquire vacuum expectation values, conformal symmetry is spontaneously broken, and all necessary mass scales—including an effective Newton’s constant and the Planck mass—emerge dynamically. This mechanism is entirely analogous to the Brout-Englert-Higgs mechanism in gauge theories and has been further developed in various contexts: Ghilencea [52] showed that the Stueckelberg mechanism in Weyl quadratic gravity generates the Planck scale while reducing Weyl geometry to Riemannian geometry at low energies; ’t Hooft [139] proposed that local conformal symmetry is a fundamental gauge symmetry of nature that is spontaneously broken through dilaton dynamics; and Salvio and Strumia [117] demonstrated in their “Agravity” framework that the Planck scale emerges via dimensional transmutation from purely dimensionless couplings.
Within this paradigm, the GUP should be understood as an effective phenomenological description of quantum gravity corrections, valid in the low-energy regime where conformal symmetry is already broken and the Planck scale has emerged. This interpretation is supported by derivations of GUP from UV-complete theories such as string theory [69] and its connection to the running Newton’s constant in asymptotically safe gravity [120]. Our approach therefore employs CWG to provide the classical gravitational background (the MK BH solution), while GUP captures the leading-order quantum corrections to matter fields propagating on this background. This effective field theory perspective ensures complete theoretical consistency: conformal symmetry governs the UV structure of the theory, while GUP-modified thermodynamics describes the IR physics where symmetry breaking has generated the relevant scales.
@@ 4 GUP Corrected Hawking Radiation of CWGBHs
To account for the profound influence of quantum gravitational effects on BH thermodynamics, one of the most promising and well-developed theoretical approaches involves the implementation of the GUP. This principle introduces a fundamental minimal length scale that emerges naturally from several candidate theories of quantum gravity, including string theory, loop quantum gravity, and non-commutative geometry [47, 141, 98]. The GUP framework provides a phenomenological window into the quantum gravitational regime by modifying the standard Heisenberg uncertainty relations through the incorporation of Planck-scale corrections.
@@
$\Gamma_{\text{GUP}}$ $\sim\exp\left(-\frac{2\pi E_{\text{GUP}}}{\kappa}\right)$
$=\exp\left(-\frac{4\pi E}{B^{\prime}(r_{h})}\left(1+\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right)\right).$ (30)
@@
$T_{\text{GUP}}$ $=\frac{B^{\prime}(r_{h})}{4\pi}\left(1+\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right)^{-1}$
$=T_{H}\left(1+\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right)^{-1},$ (31)
where $T_{H}$ represents the classical Hawking temperature derived in the previous section.
Figure 1: Quantum temperature suppression factor $\Delta T/T_{\text{classical}}$ versus horizon radius $r_{h}$ for GUP parameters $\lambda=\{0.5,1.0,1.5\}$, calculated with CWG parameters $\beta=2$, $\gamma=1$, and $k=1$. The suppression factor quantifies the fractional reduction in thermal radiation due to minimal length effects, exhibiting enhanced quantum corrections for smaller horizon radii. This computational analysis demonstrates the systematic energy scale modifications induced by the generalized uncertainty principle, with the suppression becoming dominant in regimes where $r_{h}\sim l_{p}$, establishing clear boundaries between semiclassical and quantum gravitational domains.
This result demonstrates that the emission spectrum undergoes a systematic redshift due to quantum gravitational corrections, with the magnitude of the effect being inversely related to the square of the horizon radius. As illustrated in Figure 1, the quantum suppression factor exhibits characteristic scaling behavior that becomes increasingly pronounced for smaller horizon radii, quantitatively validating the theoretical predictions of minimal length effects in gravitational thermodynamics.
@@
The phase transition implications of GUP corrections become manifest through the modified critical behavior, fundamentally altering the thermal energy scales at which thermodynamic instabilities emerge [2, 97, 143]. For CWGBHs, the quantum-corrected critical temperature exhibits the form
$T_{\text{crit,GUP}}=T_{\text{crit}}\left(1+\frac{\lambda^{2}l_{p}^{2}}{2r_{h}^{2}}\right)^{-1},$ (32)
where $T_{\text{crit}}$ represents the classical critical temperature determined by the heat capacity divergence condition $kr_{h}^{3}+3\beta^{2}\gamma-2\beta=0$ at the critical horizon radius $r_{\text{crit}}$. The explicit expression for the classical critical temperature follows from the surface gravity evaluation
$T_{\text{crit}}$ $=\frac{B^{\prime}(r_{\text{crit}})}{4\pi}$
$=\frac{3\beta^{2}\gamma}{4\pi r_{\text{crit}}^{2}}+\frac{\beta}{2\pi r_{\text{crit}}^{2}}+\frac{\gamma}{4\pi}+\frac{kr_{\text{crit}}}{2\pi},$ (33)
characterizing the thermal energy boundary separating thermodynamically stable and unstable phases in the absence of quantum gravitational corrections.
Figure 2: GUP-modified critical temperature as a function of event horizon radius $r_{h}$ for various quantum correction parameters $\lambda=\{0,0.5,1.0,1.5,2.0\}$, with fixed CWG parameters $\beta=2$, $\gamma=1$, and $k=1$. The classical limit ($\lambda=0$) exhibits the standard inverse-square scaling behavior, while increasing GUP parameters systematically suppress the critical temperature across all horizon scales. The quantum suppression becomes increasingly pronounced for smaller horizon radii, demonstrating the enhanced significance of minimal length effects in the near-Planckian regime where $r_{h}\to l_{p}$.
The GUP parameter $\lambda$ introduces a characteristic quantum length scale that systematically suppresses the critical temperature through minimal length effects, thereby modifying phase boundaries and shifting the stability regions identified through heat capacity analysis toward lower energy scales [106, 50, 110]. As demonstrated in Figure 2, the quantum-corrected critical temperature exhibits systematic suppression that becomes increasingly pronounced for smaller horizon radii, reflecting the enhanced significance of minimal length corrections in the near-Planckian regime.
These quantum corrections demonstrate systematic alterations to thermodynamic phase structures in the near-Planckian regime, where the quantum suppression factor becomes increasingly pronounced as $r_{h}\to l_{p}$, establishing domains where conventional semiclassical descriptions become inadequate for capturing the full spectrum of gravitational phenomena [29, 61].
5 QC Thermodynamics of CWGBHs
@@
To provide a more complete theoretical framework, quantum statistical mechanics offers a sophisticated approach through systematic counting of horizon microstates [55, 71]. In canonical ensembles with fixed energy and particle number, quantum fluctuations give rise to subleading modifications to the classical entropy expressions. These corrections effectively capture the statistical uncertainties arising from quantum degrees of freedom that are entirely absent in semiclassical treatments, providing crucial insights into the quantum nature of gravitational thermodynamics [35, 58, 130].
@@
We note that alternative quantum corrections to BH entropy have been proposed in the literature, most notably the logarithmic form $S=S_{0}-\frac{3}{2}\ln S_{0}+\ldots$, which arises from one-loop quantum fluctuations around the saddle-point geometry in the Euclidean path integral [38, 67, 121]. While logarithmic corrections dominate perturbative quantum effects for large BHs, the exponential form $S=S_{0}+e^{-S_{0}}$ captures non-perturbative effects associated with the discrete nature of the BH microstate spectrum [32]. Since our analysis focuses on the near-Planckian domain where $S_{0}\sim\mathcal{O}(1)$ and GUP corrections are most operative, the exponential correction provides the physically appropriate non-perturbative completion. For large BHs ($S_{0}\gg 1$), the exponential term $e^{-S_{0}}\to 0$ and the standard Bekenstein-Hawking entropy is recovered.
Regarding the choice of $S_{0}=S_{BH}=\pi r_{h}^{2}$ as the base entropy: in higher-derivative theories such as CWG, the Wald entropy [149, 62] generically differs from the Bekenstein-Hawking area law. However, for the MK vacuum solution in the symmetry-broken phase—where the effective Newton constant $G_{\rm eff}$ emerges dynamically [85, 53]—the Wald entropy evaluated on the static, spherically symmetric horizon reduces to the standard area law $S_{\rm Wald}=\pi r_{h}^{2}$ (in units $G_{\rm eff}=1$[76]. This equivalence holds because the Weyl-squared contributions to the Wald integrand produce only area-proportional terms for Einstein space solutions on a two-sphere horizon cross-section [101]. Therefore, $S_{0}=S_{BH}=\pi r_{h}^{2}$ is the appropriate starting point for the exponential correction scheme applied to the MK metric.
It is important to emphasize that the thermodynamic quantities presented in this section are computed as parametric functions of the horizon radius $r_{h}$ and the CWG parameters $(\beta,\gamma,k)$. This parametric approach is standard in BH thermodynamics and reveals the functional dependence of thermodynamic potentials on horizon geometry. For numerical calculations requiring specific horizon values, one should employ the self-consistent approach where $r_{h}$ is determined from $B(r_{h})=0$, as detailed in Section 3. Table 2 presents explicit numerical values of the Bekenstein-Hawking entropy $S_{BH}=\pi r_{h}^{2}$ and Hawking temperature $T_{H}$ computed using self-consistent horizon radii for various CWG parameter combinations.
$\beta$ $\gamma$ $k$ $r_{h}$ $S_{BH}$ $T_{H}$
0.5 0.00 0.00 1.0000 3.1416 0.079577
1.0 0.00 0.00 2.0000 12.5664 0.039789
2.0 0.00 0.00 4.0000 50.2655 0.019894
0.5 0.05 0.00 0.9878 3.0654 0.082476
1.0 0.05 0.00 1.9523 11.9738 0.042605
2.0 0.05 0.00 3.8167 45.7631 0.022553
0.5 0.10 0.00 0.9761 2.9934 0.085210
1.0 0.10 0.00 1.9083 11.4408 0.045106
2.0 0.10 0.00 3.6569 42.0112 0.024620
0.5 0.00 0.05 0.9563 2.8729 0.094631
1.0 0.00 0.05 1.7377 9.4859 0.066538
2.0 0.00 0.05 2.8466 25.4574 0.061934
0.5 0.05 0.05 0.9463 2.8134 0.097038
1.0 0.05 0.05 1.7102 9.1884 0.067923
2.0 0.05 0.05 2.7769 24.2250 0.061164
Table 2: Bekenstein-Hawking entropy $S_{BH}=\pi r_{h}^{2}$ and Hawking temperature $T_{H}$ for CWGBHs with self-consistent horizon radii $r_{h}$ determined from $B(r_{h})=0$. The first three rows correspond to the Schwarzschild limit ($\gamma=0$, $k=0$), where $r_{h}=2\beta$ and $S_{BH}=4\pi\beta^{2}$ are exactly recovered. The CWG parameters systematically reduce the horizon radius and entropy while increasing the temperature, consistent with the thermodynamic principle that smaller BHs radiate more intensely.
@@
$E=\int T_{H}\,dS,$ (35)
where $T_{H}$ denotes the Hawking temperature derived in Eq. (25). For macroscopic BHs with $S_{0}\gg 1$, the exponential correction $e^{-S_{0}}$ is negligible (e.g., $e^{-\pi r_{h}^{2}}\approx 4.3\times 10^{-2}$ at $r_{h}=1$ and $\approx 3.5\times 10^{-6}$ at $r_{h}=2$), so the integration measure reduces to $dS\approx dS_{0}=2\pi r_{h}\,dr_{h}$, with the expansion parameter being $e^{-S_{0}}$. The $\approx$” signs in the following expressions indicate that terms of order $\mathcal{O}(e^{-\pi r_{h}^{2}})$ have been dropped. Substituting the QC entropy expression and performing the integration yields
$E\approx\left(\beta-\tfrac{3}{2}\beta^{2}\gamma\right)\ln r_{h}+\frac{\gamma\,r_{h}^{2}}{4}+\frac{k\,r_{h}^{3}}{3}.$ (36)
This expression is defined up to an additive constant of integration [66]. In the Schwarzschild limit ($\gamma=0$, $k=0$), it reduces to $E=\beta\ln r_{h}$, which has the correct dimension $\text{mass}]\times[\text{dimensionless}]$. The first-law consistency $dE/dr_{h}=T_{H}\cdot dS_{0}/dr_{h}$ holds identically by construction.
@@
Figure 3 presents a filled contour plot illustrating the behavior of the internal energy $E$ as a function of both the event horizon radius $r_{h}$ and the scale-dependent parameter $\gamma$, with fixed parameter values $\beta=2$ and $k=1$. The visualization clearly demonstrates that the energy increases monotonically with $r_{h}$, consistent with the thermodynamic interpretation of BH mass. The contour lines, which appear nearly vertical across the parameter space, indicate that $E$ depends predominantly on $r_{h}$ with only weak sensitivity to $\gamma$. The colorbar scale ranges from approximately 0 to $55$ units. The contour structure reveals that $\gamma$ exerts a negligible influence on $E$ at larger horizon scales, but becomes increasingly important in the small $r_{h}$ regime. This behavior indicates that scale-dependent geometric corrections subtly modify the BH energy structure near the Planckian regime, suggesting that quantum corrections encoded in $\gamma$ become critically relevant in the small-scale limit and may significantly influence evaporation dynamics or remnant formation scenarios.
Figure 3: Filled contour plot of the internal energy $E$ as a function of the event horizon radius $r_{h}$ and the scale-dependent parameter $\gamma$, with fixed values $\beta=2$ and $k=1$. The contour lines with colorbar quantitatively demonstrate the rapid increase of energy with $r_{h}$, while the near-vertical contour structure confirms the minimal influence of $\gamma$ across most of the parameter space, highlighting the dominant role of horizon size in determining the BH mass-energy content.
@@
The Helmholtz free energy, which provides crucial information about the system’s equilibrium properties, is defined through the standard thermodynamic relation [129]
$F=E-T_{H}S_{0},$ (37)
yielding upon substitution
$F\approx\left(\beta-\tfrac{3}{2}\beta^{2}\gamma\right)\ln r_{h}-\frac{k\,r_{h}^{3}}{6}+\frac{3\beta^{2}\gamma}{4}-\frac{\beta}{2}.$ (38)
Figure 4 displays the Helmholtz free energy behavior, revealing the complex interplay between geometric parameters and thermodynamic stability. The filled contour representation with colorbar shows $F$ taking predominantly positive values (yellow regions) at small $r_{h}$ and transitioning to increasingly negative values (purple regions) as $r_{h}$ increases, with the colorbar indicating a range from approximately 0 to 30-30 units. The contour structure demonstrates that larger BHs possess more negative free energy, indicating enhanced thermodynamic stability. The slight curvature of contour lines at small $r_{h}$ reveals the region where $\gamma$-dependent corrections become non-negligible.
Figure 4: Filled contour plot of the Helmholtz free energy $F$ displaying the QC thermodynamic landscape for CWGBHs, with fixed parameters $\beta=2$ and $k=1$. The colorbar quantifies the transition from positive (yellow, small $r_{h}$) to negative (purple, large $r_{h}$) free energy values, with contour lines illustrating the complex parameter dependence that governs equilibrium stability conditions. The predominantly vertical contour structure at large $r_{h}$ indicates $r_{h}$-dominated behavior, while deviations at small $r_{h}$ reveal $\gamma$-dependent quantum corrections.
@@
The QC pressure, fundamental for understanding the mechanical properties of the BH system, follows from the thermodynamic relation [58]
@@
$P\approx\frac{k\,r_{h}^{3}+3\beta^{2}\gamma-2\beta}{8\pi\,r_{h}^{3}}.$ (40)
In the Schwarzschild limit ($\gamma=0$, $k=0$), this reduces to $P=-\beta/(4\pi r_{h}^{3})$.
Figure 5 presents the QC pressure as a filled contour plot, revealing a rich two-dimensional structure that distinguishes it from the other thermodynamic potentials. The colorbar scale ranges from approximately 0.6-0.6 to $1.4$ units, with the contour lines exhibiting significant curvature throughout the $(r_{h},\gamma)$ parameter space. Notably, the pressure shows strong dependence on both parameters: at fixed $r_{h}$, increasing $\gamma$ systematically increases the pressure, while at fixed $\gamma$, the pressure grows with $r_{h}$. The region of negative pressure (dark purple) at small $r_{h}$ and small $\gamma$ corresponds to configurations where quantum corrections and geometric effects compete, potentially indicating exotic thermodynamic behavior characteristic of the CWG framework.
Figure 5: Filled contour plot of the QC pressure $P$ as a function of horizon parameters $r_{h}$ and $\gamma$, with fixed $\beta=2$ and $k=1$. The colorbar quantifies pressure values ranging from negative (purple) to positive (yellow), while the curved contour structure illustrates the non-trivial dependence on both parameters. The mechanical response properties of CWGBHs exhibit significantly richer structure than the energy-type potentials, with the $\gamma$ parameter playing a more pronounced role in determining pressure characteristics.
@@
resulting in the expression
$H$ $\approx\left(\beta-\tfrac{3}{2}\beta^{2}\gamma\right)\ln r_{h}+\frac{\gamma\,r_{h}^{2}}{4}$
$\quad+\frac{k\,r_{h}^{3}}{2}+\frac{\beta^{2}\gamma}{2}-\frac{\beta}{3}.$ (42)
Figure 6 displays the enthalpy distribution as a filled contour plot, demonstrating behavior qualitatively similar to the internal energy $E$. The colorbar scale extends from approximately 0 to $85$ units. The nearly vertical contour lines confirm that $H$ is predominantly determined by $r_{h}$, with $\gamma$ producing only minor modifications. This $r_{h}$-dominated behavior reflects the fact that both $E$ and $PV$ scale strongly with horizon radius, making the enthalpy an excellent proxy for the total thermodynamic energy content of the BH system.
Figure 6: Filled contour plot of the enthalpy $H$ demonstrating the total energy content of CWGBHs under constant pressure conditions, with fixed parameters $\beta=2$ and $k=1$. The colorbar quantifies enthalpy values, while the near-vertical contour structure reveals the quantum-enhanced thermodynamic behavior dominated by horizon radius dependence.
@@
$G\approx-\frac{1}{12}\left(6\ln r_{h}-5\right)\left(3\beta\gamma-2\right)\beta.$ (43)
Figure 7 presents the Gibbs free energy as a filled contour plot, with the colorbar indicating values ranging from approximately 3-3 to +5+5 units. Unlike the original polynomial expression, the corrected $G$ exhibits sign changes—specifically at $r_{h}=e^{5/6}\approx 2.30$—indicating thermodynamic phase boundaries. The contour structure shows predominantly vertical lines with slight curvature, indicating that while $r_{h}$ remains the dominant parameter, the $\gamma$ dependence becomes more pronounced at larger horizon radii.
Figure 7: Filled contour plot of the Gibbs free energy $G$ characterizing the global thermodynamic stability landscape and phase transition structure of QC CWGBHs, with fixed parameters $\beta=2$ and $k=1$. The colorbar quantifies $G$ values ranging from 3-3 to +5+5 units. The sign change of $G$ at $r_{h}\approx e^{5/6}$ indicates a thermodynamic phase boundary, with the $\gamma$-dependent modifications becoming more apparent at larger horizon scales.
@@
leading to the result
$C\approx\frac{\pi\left(2k\,r_{h}^{3}-3\beta^{2}\gamma+\gamma r_{h}^{2}+2\beta\right)r_{h}^{2}}{k\,r_{h}^{3}+3\beta^{2}\gamma-2\beta}.$ (45)
Note the prefactor $\pi r_{h}^{2}$ (rather than $\pi^{2}r_{h}^{4}$), which ensures the correct Schwarzschild limit $C=-\pi r_{h}^{2}=-S_{BH}$ when $\gamma=0$ and $k=0$.